OpenStudy (anonymous):

calculate the derivative with respect to x of y + 1/y=x^2+x

OpenStudy (anonymous):

I know you have to take the derivative of both sides

OpenStudy (anonymous):

not quite where to go from here

OpenStudy (abb0t):

simply take the derivatice of y as you would for x, except add a y', since y is a function of x, you must include that. for example: \[\frac{ d }{ dx }(y^2) = 2yy' = 2y \frac{ dy }{ dx }\]

OpenStudy (abb0t):

Then, you can factor out a dy/dx and solve for dy/dx

OpenStudy (anonymous):

k. can u go throgh the steps please. Y^2 ?

OpenStudy (abb0t):

\[y' - \frac{ 1 }{ y^2 }y'\] is the left side. Now, do the right side. group like terms (remember from algebra). factor out and solve for y'. \[(ax+ay) = a(x+y)\]

OpenStudy (abb0t):

for \[\frac{ d }{ dx }(a^n) = na^{n-1} \] remember

OpenStudy (anonymous):

why\[y + \frac{ 1 }{ y }= x^2 + x\]

OpenStudy (abb0t):

\[\frac{ 1 }{ y } = y^{-1}\]

OpenStudy (anonymous):

want to make sure the equations was inputted correctly

OpenStudy (abb0t):

So, take the derivative as you would for the whole thing.

OpenStudy (abb0t):

except whenever u take derivative of "y" in this, you multiply it by y'

OpenStudy (abb0t):

the derivative is to tell you that it's just a derivative with respect to x. Kind of like chain rule.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

is the first step d / dx (y) + d / dx (1/y) = d / dx ( x^2 + x )

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 2x+1 }{1-\frac{ 1 }{ y^2 } }\]

OpenStudy (anonymous):

can you show me the first step; just getting used to using the equation button

OpenStudy (anonymous):

i know you have to take the derivative of both sides d/ dx (y) + d/dx ( 1 / y ) = d / dx (x^2 + x)

OpenStudy (anonymous):

then do I write in the y' on each side

OpenStudy (anonymous):

okay let me write it in full

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

okay let e solve it on my book then i am gonna attach it as a pic coz it takes time having to do it here

OpenStudy (anonymous):

yes it does; thanks

OpenStudy (anonymous):

I am trying another while waiting for the pic

OpenStudy (anonymous):

\[\sin \left( x+y \right)=\left( x+\cos y \right)\]

OpenStudy (anonymous):

Will I see the pic in this screen

OpenStudy (anonymous):

yes i guess

OpenStudy (anonymous):

did you already post the answer in full

OpenStudy (anonymous):

can u see it?

OpenStudy (anonymous):

yes. I will try to open it

OpenStudy (anonymous):

where did the 1 in the 1 - 1 / y squared come from

OpenStudy (anonymous):

do I have to write the y squared beside the (2x + 1)

OpenStudy (anonymous):

its like x-xy then you can write x(1-y) so thats how i got the 1

OpenStudy (anonymous):

you can simplify it according to how you understand it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can you go through another one; I already posted it: sin ( x + y) = (x + cos y)

OpenStudy (anonymous):

I started with \[\frac{ d }{ dx }\sin x)+ \]

OpenStudy (anonymous):

d / dx (sin x) + d/dx(sin y) = d/dx (x) + d/dx (cos y)

OpenStudy (anonymous):

okay i`m on it

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

\[\frac{ dy }{ dy }=\frac{ 1-\cos(x) }{ 1+\sin(y) }\] thats what i got what did u get

OpenStudy (anonymous):

\[y`=(\cos x + y ) = 1 - \sin y)\]

OpenStudy (anonymous):

oh sorry my mistake

OpenStudy (anonymous):

I know I missed a step

OpenStudy (anonymous):

i forgot to differentiate the inner function

OpenStudy (anonymous):

the book`s answer has 1- cos(x+y) / cos ( x + y) + sin y

OpenStudy (anonymous):

not sure how to get to the books answer from what your anser and my answer

OpenStudy (anonymous):

could you verify how to differentiate the inner function

OpenStudy (anonymous):

dy/dx=sin(x+y) Put x+y=u then,differentiate

OpenStudy (anonymous):

yeah i am on it

OpenStudy (anonymous):

Great- you are a big help!

OpenStudy (anonymous):

its a mind challenging but i never give up i am stil coming with a solution not long

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

\[\frac{ d(\sin(x+y)) }{ dx }=\cos(x+y)(1+\frac{ dy }{ dx })\] so we can take it from here

OpenStudy (anonymous):

\[\frac{ dy }{dx }=\frac{ 1-\cos(x+y) }{ \sin(y)+\cos(x+y)) }\] thats what i got check it tell me is i got it wrong even now

OpenStudy (anonymous):

that`s exactly what is in the answer key of the book

OpenStudy (anonymous):

what was your second step; before the answer

OpenStudy (anonymous):

 something like that